In an accelerator center, an experiment needs a 1.41 cm thickaluminum cylinder. Suppose that the thickness of a cylinderhas a normal distribution with a mean 1.41 cm and a standarddeviation of 0.01 cm.


1
Expert's answer
2021-10-14T18:15:36-0400

Let "X=" the thickness of a cylinder: "X\sim N(\mu, \sigma^2)."

Given "\mu=1.41\ cm, \sigma=0.01\ cm."

(a) What is the probability that a thickness is greater than 1.42 cm?


"P(X>1.42)=1-P(X\leq 1.42)"

"=1-P(Z\leq\dfrac{1.42-1.41}{0.01})=1-P(Z\leq1)"


"\approx0.158655"

(b) What thickness is exceeded by 95% of the samples?


"P(X>x)=1-P(X\leq x)"

"=1-P(Z\leq\dfrac{x-1.41}{0.01})=0.95"

"P(Z\leq\dfrac{x-1.41}{0.01})=0.05"

"\dfrac{x-1.41}{0.01}\approx-1.6449"

"x=1.41-0.16449"

"x=1.2455"

Thickness "1.2455" cm is exceeded by 95% of the samples.


(c) If the specifications require that the thickness is between 1.39 cm and 1.43 cm, what proportion of the samples meets specifications?


"P(1.39<X<1.43)=P(X<1.43)-P(X\leq 1.39)"

"=P(Z<\dfrac{1.43-1.41}{0.01})-P(Z\leq\dfrac{1.39-1.41}{0.01})"

"=P(X<2)-P(Z\leq-2)"

"=0.97725-0.02275=0.9545"

95.45 % of the samples meet specifications.